OpenStudy (anonymous):

Solve 1 + 2/x + 5/x^2 = 0

OpenStudy (anonymous):

Multiply both sides by x^2 to get rid of fractions, then solve as linear equation.

OpenStudy (anonymous):

okay I'm going to try….

OpenStudy (anonymous):

thank you but don't go anywhere in case i need help

OpenStudy (anonymous):

wait why x^2

OpenStudy (anonymous):

cause you have 5/x^2 at the end, right? So you need to get rid of that

OpenStudy (anonymous):

uh um im confused….

OpenStudy (anonymous):

multiplying it by x^2 will also convert 2/x into 2x (2/x * x^2 = 2x)

OpenStudy (anonymous):

ok so if you multiply this entire equation by x^2 here is what will happen. Your right side will remain 0, since 0*x^2 = 0 Your left side will become this: x^2 + 2x + 5 So by doing that you get equation: x^2 + 2x + 5 = 0 which is your basic quadratic equation or whatever they called and could be solved using formula

OpenStudy (anonymous):

okay so is it X^2 + 2x + 5 = 0

OpenStudy (anonymous):

oh haha okay thank you but how did you know we could use X^2 like why not just x as the greatest common factor?

OpenStudy (anonymous):

Well you need to get rid of fraction and as a rule thumb the easiest way to find the factor is to look at the fraction with highest denominator, right? :)

OpenStudy (anonymous):

oh yeah okay….fractions with variables scare, i was probably over thinking.

OpenStudy (anonymous):

THATN YOU!!!!

OpenStudy (anonymous):

np, glad to be of help.